# Xingzhi Zhou1,3, Xin Dong2, Chunhao Li1, Yuning Bai4, Yulong Xu5, Ka Chun Cheung3, Simon See6, Xinpeng Song2 ✉Runshun Zhang3, ✉Xuezhong Zhou2, and ✉Nevin L. Zhang1

### v1

@@ III-D Experimental setup
We implement the fine-tuning process with the Transformers package from Hugging Face [^2] https://huggingface.co/docs/transformers. The number of training epochs is set to 10. A cosine scheduler is adopted, with the rank of LoRA being 16 and the alpha of LoRA at 32. We consider the number of herb permutations, K, as 20 and 50, respectively. The default learning rate is set at 1e-3. We leverage the gradient accumulation technique to increase the batch size, with 8 accumulation steps and a batch size of 16. We employ two foundation models Chinese-Alpaca-Plus-7B  [^3] https://github.com/ymcui/Chinese-LLaMA-Alpaca/tree/main and ShenNong  [^4] https://huggingface.co/michaelwzhu/ShenNong-TCM-LLM. Chinese-Alpaca-Plus-7B is a variant of LLaMA [25] with continual pre-training and supervised fine-tuning on Chinese corpus, denoted as LLaMA+ for simplicity. ShenNong is a further refinement of LLaMA+, fine-tuned with TCM instruction datasets. We use an 8-V100 GPU machine for fine-tuning and the running time for TCM-FTP ($K=50$) is 146 hours. During the inference stage, We employ top-k and top-p combinations for decoding, setting the top-k to 50 and the top-p to 0.7, with a default temperature of 0.95.
@@ III-E Evaluation Metrics
For the proposed TCM-FTP, we evaluate it from both quantitative and qualitative perspectives, including the following evaluation metrics.
Quantitative Evaluation. For evaluation metrics, we use precision, recall, and F1-score as herb prediction metrics and designed NMSE for herb dosage evaluation.
  • Precision: The proportion of correctly predicted herbs out of all predicted herbs, reflecting the accuracy of positive predictions.
  • Recall: The proportion of correctly predicted herbs out of all herbs in the ground truth, indicating the model’s ability to identify all relevant items.
  • F1-score: The harmonic mean of precision and recall, giving a balanced measure of overall accuracy.
  • Normalized mean square error (NMSE): To tackle the problem above, we design this metric to evaluate the accuracy of predicted dosage by normalizing the squared differences using the original weights. For a given prescription $p$ and a predicted prescription and $\hat{p}$, we suppose they are composed by a set of pairs of a herb and a dosage, $p=\{(h,w)\}$, where $h,w$ refer to herb and dosage respectively, and $\hat{p}=\{(h^{\prime},w^{\prime})\}$ indicates the generated result from model. The calculation of NMSE is as follows,
    $NMSE=\frac{1}{Z}\sum_{\begin{subarray}{c}(h,w)\in p,\\ \left(h^{\prime},w^{\prime}\right)\in\hat{p}\end{subarray}}\mathbf{1}\left[h=h^{\prime}\right]\left(\frac{w^{\prime}-w}{w}\right)^{2},$ (9)
    where $Z$ is the number of correctly predicted herbs, and $\mathbf{1}[\cdot]$ is the indicator function. For evaluation, the average dosage of each herb from the training data is used as a baseline for dosage predictions. For herbs unseen in the training data, the dosage is predicted as the average dosage of all known herbs. This approach is referred to as NMSEbase in the NMSE calculations.
@@ IV-A Overall Results
MLC models & topic models MLKNN [42] 0.5365 0.4626 0.4968 - -
LinkLDA [37] 0.5267 0.4572 0.4895 - -
LinkPLSALDA [39] 0.5311 0.4614 0.4938 - -
[…]
@@
The overall comparison results between TCM-FTP and the baselines are presented in Table II. The proposed TCM-FTP outperforms all baseline models in the herb prediction task on DigestDS, achieving an F1-score of 0.8016 using LLaMA+ as the foundation model and 0.8031 using ShenNong. This highlights the superior capability of TCM-FTP in modeling the intricate relationships between symptoms and herbs in prescriptions. Unlike previous approaches, TCM-FTP also includes herb dosage prediction, which is crucial in TCM due to the significant impact of dosage combinations. As shown in Table II, TCM-FTP achieves a much lower NMSE compared to the baseline using average statistics (0.0604 for TCM-FTP versus 0.0754 for Mengzi (T5-base) [45]). This enhances the practicality of TCM-FTP in clinical TCM prescription generation.
Our proposed TCM-FTP shows significant advantages over various baseline models across all aspects.
  • Compared to the pre-trained language model Mengzi (T5-base) [45], TCM-FTP significantly enhances prediction performance with improved precision (from 0.7332 to 0.7951), recall (from 0.7474 to 0.8113), and F1-score (from 0.7403 to 0.8031). Performances on GPT-3.5 and GPT-4.0 [24] were notably poor, indicating the limitations of general llms in tasks requiring specialized knowledge.
@@
  • In comparison to other TCM prescription prediction models, TCM-FTP outperforms PTM [5], which leads among these baselines, followed by TCMPR [43] in accuracy, with KDHR [44] and PresRecST [12] showing poorer performance. PTM, a topic model-based approach, is computationally complex, while TCMPR, KDHR, and PresRecST are graph-based models sensitive to graph-related factors. TCM-FTP excels these approaches in generating accurate herb recommendations due to advanced language modeling capability.
  • Compared to multi-label classification (MLC) and topic model approaches, our model demonstrates superior performance. MLKNN [42] and two topic model methods perform similarly, whereas MLC and topic model-based approaches achieve lower performance due to computational intensity and limited consideration of inter-herb relationships. This underscores the proposed TCM-FTP proves more suitable for real-world herb prediction tasks, highlighting its enhanced performance over various baseline methods.
In summary, our approach exhibits significant advantages over various baseline methods and is better suited for real-world herb recommendation tasks. In addition, TCM knowledge embedded in foundation models enhances the fine-tuning process. Epoch-wise performance, shown in Fig. 3(a) for both loss and F1-score, indicates that although TCM-FTP (ShenNong, K=50) shows similar loss changes to TCM-FTP (LLaMA+, K=50), the F1-score line for TCM-FTP (ShenNong, K=50) consistency surpasses that of TCM-FTP (LLaMA+, K=50). This suggests that TCM knowledge significantly aids the fine-tuning process with the ShenNong model. Additionally, with a lower number of permutations K=20, TCM-FTP (ShenNong) achieves an F1-score of 0.8008, compared to 0.7652 for TCM-FTP (LLaMA+), highlighting the advantage of having TCM knowledge embedded in the ShenNong model for modeling symptom-herb relationships.
IV-B Parameter Analysis
The number of permutation The herb permutation introduces the order-agnostic property to enhance the prediction performance. We present the results of TCM-FTP with varying numbers of herb permutations $K$ in Table III. As $K$ increases from 0 to 50, the f1-score improves gradually from 0.4885 to 0.8031, and the NMSE decreases from 0.1683 to 0.0604. This suggests the importance of the order-agnostic property as an effective inductive bias for prescription prediction.
TABLE III: The impact of herb permutations TCM-FTP is evaluated with different numbers of herb permutations, K.
K Precision Recall F1-score NMSE NMSEbase
0 0.4640 0.5156 0.4885 0.1683 0.1509
1 0.5572 0.5765 0.5667 0.1269 0.1410
10 0.7748 0.7902 0.7824 0.0608 0.1391
20 0.7919 0.8100 0.8008 0.0607 0.1441
50 0.7951 0.8113 0.8031 0.0604 0.1431
Learning rates We validate the impact of learning rates by conducting experiments with K=10 and varying learning rates, as shown in Table IV. Higher learning rates correlate with improved precision, recall, and F1-score. This suggests that lower learning rates lead to underfitting, likely due to significant disparities between the pre-trained model parameters and the target parameters. Our objective is to fine-tune a task-specific generator without retaining the broad language capabilities of large language models (LLMs), necessitating substantial adjustments to achieve superior performance. However, excessively large learning rates cause model corruption, such as a learning rate of 0.0025.
TABLE IV: Impact of learning rates TCM-FTP (ShenNong, K=10) is fine-tuned with different learning rates.
lr Precision Recall F1-score NMSE NMSEbase
5e-5 0.5766 0.5894 0.5830 0.1057 0.1238
1e-4 0.6395 0.6615 0.6503 0.0869 0.1282
5e-4 0.7571 0.7734 0.7652 0.0693 0.1475
7.5e-4 0.7707 0.7832 0.7769 0.0824 0.1394
1e-3 0.7748 0.7902 0.7825 0.0608 0.1391
2.5e-3 0.0003 0.0000 0.0000 0.0178 0.0400
Decoding paramters We investigate the robustness of decoding parameters by varying top-p, top-k, and temperature. Initially, in Fig. 3(b), we set top-p to 0.7 and chart the f1-score across various top-k and temperature settings. The results show that the model’s performance is stable across different temperatures and top-k values, with only a slight decline when the temperature exceeds 0.95. Next, in Fig. 3(c), with top-k fixed at 50, we graph the f1-score for different top-p and temperature values, observing that lower temperatures are more stable, leading to higher probabilities for the top predicted tokens, thus enhancing prediction accuracy. Lastly, in Fig. 3(d), with the temperature set at 0.95, we map the f1-score for various top-p and top-k combinations, noting that lower top-k values offer more stability in the face of top-p variations.
(a) Loss and F1-score versus epoches with TCM-FTP.
(b) Top-k versus temperature under a fixed top-p
(c) Top-p versus temperature under a fixed top-k
(d) Top-p versus top-k under a fixed temperature
Figure 3: Epoch-wise performance and decoding parameter analysis.
IV-C Expert Qualitative Evaluation
@@
Since conducting expert qualitative evaluation (EQE) is time-consuming and labor-intensive for doctors, this study randomly selected 20 data points from the test set for evaluation. As shown in Table V, the TCM-FTP model outperformed other baseline models across all metrics, achieving an average SHE score of 4.08 (standard deviation 0.5628) and an average HC score of 4.03 (standard deviation 0.5404). This indicates that TCM-FTP can generate effective prescriptions that adhere to TCM compatibility principles. In contrast, GPT-3.5 and GPT-4.0 [24] scored below 3 for both SHE and HC, with total scores of 5.75 and 5.89, respectively, significantly lower than TCM-FTP and TCMPR [43]. This demonstrates that GPT-3.5 and GPT-4.0 perform poorly in generating prescriptions that adhere to TCM principles. TCMPR performed relatively well, with SHE and HC scores of 3.56 (standard deviation 0.6715) and 3.54 (standard deviation 0.6578), respectively, but still fell short of TCM-FTP. Overall, TCM-FTP excelled in prescription generation, receiving higher approval from doctors and showing less score variability, indicating better stability and robustness.
@@
TABLE V: Comparative Results of expert evaluation. The table records the mean scores (standard deviation) for all evaluated samples.
@@
IV-D Case Study
In order to visually showcase the predictive performance and capabilities of our TCM-FTP, we obtain the predicted herb results formed by each model on the test set and had them evaluated by TCM experts. Fig. 4 illustrates the results for two cases, including the input (chief complaints, present medical history), output (actual prescriptions provided by the doctor), and the predicted prescription results from TCM-FTP (ShenNong) and Mengzi (T5 base) [45] models, as well as GPT-3.5 and GPT-4.0 [24] predictions. Herb names/dosage weights marked in red indicate that the model’s prediction match the actual prescription provided by the doctor. Additionally, professional evaluations of each model are presented, ranking them using the “>” symbol (with models ranked higher indicating better performance).
Figure 4: Two Specific Case Analyses. This figure presents two specific test data, including inputs, outputs, predictions from each model, and evaluations from experts. Herb names/dosage weights marked in red indicate the results consistent with the expert-prescribed outcomes.
The results show that the TCM-FTP (ShenNong) models achieve strong prediction performance. The herb label predictions closely match the doctor’s prescriptions, indicating effective data fitting after fine-tuning. Additionally, the dosage predictions align well with the actual dosages provided by the doctor, matching clinically relevant herb dosages. In contrast, GPT-3.5 and GPT-4.0 were also tested on the same cases but produced comparatively poor predictions with significant deviations from the actual results. This likely stems from their lack of specialized TCM training, hindering their performance.
In summary, the proposed TCM-FTP (ShenNong) outperforms general models. This highlights the importance of fine-tuning large language models on high-quality, domain-specific data to fully leverage their capabilities in specialized fields.
@@ V CONCLUSION
To deal with the lack of high-quality datasets and to improve the performance in Traditional Chinese Medicine (TCM) prescription predictions, we build a TCM prescription dataset DigestDS from clinical records and propose TCM-FTP to fine-tune large language models to predict herbs with the corresponding dosages. We collect DigestDS from practical clinical records by focusing on digestion disorder diseases. TCM-FTP employs a low-rank adaptation for computational and storage efficiency and adapts a data augmentation by randomly permutating the order of herbs in prescriptions. The experimental results reveal the remarkable effectiveness of TCM-FTP, surpassing previous methods by large margins in precision, recall, and F1-score. Additionally, our method achieved the best results in NMSE, effectively forming accurate herb and dosage predictions. In future work, we will continue to incorporate domain knowledge into model construction to further enhance performance, aiming to develop a practically usable prescription prediction model.

### v2

@@ III-D Experimental setup
We implement the fine-tuning process with the Transformers package from Hugging Face [^2] https://huggingface.co/docs/transformers. The number of training epochs is set to 10. We adopt a cosine scheduler, with the LoRA rank set to 16 and the LoRA alpha to 32. We consider the number of herb permutations, K, as 20 and 50, respectively. The default learning rate is set to $1\times 10^{-3}$. We leverage the gradient accumulation technique to increase the batch size, with 8 accumulation steps and a batch size of 16. We employ two foundation models Chinese-Alpaca-Plus-7B  [^3] https://github.com/ymcui/Chinese-LLaMA-Alpaca/tree/main and ShenNong  [^4] https://huggingface.co/michaelwzhu/ShenNong-TCM-LLM. Chinese-Alpaca-Plus-7B is a variant of LLaMA [17] with continual pre-training and supervised fine-tuning on Chinese corpus, denoted as LLaMA+ for simplicity. ShenNong is a further refinement of LLaMA+, fine-tuned with TCM instruction datasets. We use an 8-V100 GPU machine for fine-tuning and the running time for TCM-FTP ($K=50$) is 146 hours. During the inference stage, We employ top-k and top-p combinations for decoding, setting the top-k to 50 and the top-p to 0.7, with a default temperature of 0.95.
@@ III-E Evaluation Metrics
Quantitative Evaluation.
  • Precision, recall, and F1-score: Precision is the proportion of correct herb predictions, recall is the proportion within true herbs, and F1-score is their harmonic mean.
  • Normalized mean square error (NMSE): We design a metric to evaluate the accuracy of predicted dosage by normalizing the squared differences using the original weights. For a given prescription $p$ and a predicted prescription $\hat{p}$, we suppose they are composed of a set of pairs of a herb and a dosage, $p=\{(h,w)\}$, where $h,w$ refer to herb and dosage respectively, and $\hat{p}=\{(h^{\prime},w^{\prime})\}$ indicates the generated result from model. The NMSE is defined as: $NMSE=\frac{1}{Z}\sum_{\begin{subarray}{c}(h,w)\in p,\\ \left(h^{\prime},w^{\prime}\right)\in\hat{p}\end{subarray}}\mathbf{1}\left[h=h^{\prime}\right]\left(\frac{w^{\prime}-w}{w}\right)^{2},$ where $Z$ is the number of correctly predicted herbs, and $\mathbf{1}[\cdot]$ is the indicator function. For baseline, we use the average dosage of each herb from the training data for dosage predictions.
@@ IV-A Overall Results
TCM prescription prediction models PTM [3] 0.5372 0.5777 0.5567 - -
@@
As shown in Table II, the proposed TCM-FTP outperforms all baseline models in the herb prediction task on DigestDS, achieving an F1-score of 0.8016 using LLaMA+ as the foundation model and 0.8031 using ShenNong. This highlights the superior capability of TCM-FTP in modeling the intricate relationships between symptoms and herbs in prescriptions. Unlike previous approaches, TCM-FTP also includes herb dosage prediction, which is crucial in TCM due to the significant impact of dosage combinations. As shown in Table II, TCM-FTP achieves a much lower NMSE compared to the baseline using average statistics (0.0604 for TCM-FTP versus 0.0754 for Mengzi (T5-base) [26]). This enhances the practicality of TCM-FTP.
@@
Compared to the pre-trained language model Mengzi (T5-base) [26], TCM-FTP significantly enhances prediction performance with improved precision (from 0.7332 to 0.7951), recall (from 0.7474 to 0.8113), and F1-score (from 0.7403 to 0.8031). Performances on GPT-3.5 and GPT-4.0 [16] were notably poor, indicating the limitations of general LLMs in tasks requiring specialized knowledge. In comparison to other TCM prescription prediction models, TCM-FTP outperforms PTM [3], which leads among these baselines, followed by TCMPR [24] in accuracy, with KDHR [25] and PresRecST [10] showing poorer performance. PTM, a topic model-based approach, is computationally complex, while TCMPR, KDHR, and PresRecST are graph-based models sensitive to graph-related factors. TCM-FTP excels these approaches in herb recommendations due to advanced language modeling capability.
IV-B Expert Qualitative Evaluation
@@
Since conducting expert qualitative evaluation (EQE) is time-consuming and labor-intensive for doctors, this study randomly selected 20 data points from the test set for evaluation. As shown in Table III, the TCM-FTP model outperformed other baseline models across all metrics, achieving an average SHE score of 4.08 (standard deviation 0.5628) and an average HC score of 4.03 (standard deviation 0.5404). This indicates that TCM-FTP can generate effective prescriptions that adhere to TCM compatibility principles. In contrast, GPT-3.5 and GPT-4.0 [16] scored below 3 for both SHE and HC, with total scores of 5.75 and 5.89, respectively, significantly lower than TCM-FTP and TCMPR [24]. This demonstrates that GPT-3.5 and GPT-4.0 perform poorly in generating prescriptions that adhere to TCM principles. TCMPR performed relatively well, with SHE and HC scores of 3.56 (standard deviation 0.6715) and 3.54 (standard deviation 0.6578), respectively, but still fell short of TCM-FTP. Overall, TCM-FTP excelled in prescription generation, receiving higher approval from doctors and showing less score variability, indicating better stability and robustness.
@@
TABLE III: Comparative Results of expert evaluation. The table records the mean scores (standard deviation).
@@
Figure 3: Two Specific Case Analyses. We present two specific test cases, including inputs, outputs, and predictions of the models with evaluations of experts. Herb names/dosage weights are marked in red for corrected ones.
IV-C Case Study
In order to visually showcase the predictive performance and capabilities of our TCM-FTP, we obtain the predicted herb results formed by each model on the test set and had them evaluated by TCM experts. Fig. 3 illustrates the results for two cases, including the input (chief complaints, present medical history), output (actual prescriptions provided by the doctor), and the predicted prescription results from TCM-FTP (ShenNong) and Mengzi (T5 base) [26] models, as well as GPT-3.5 and GPT-4.0 [16] predictions. Herb names/dosage weights marked in red indicate that the model’s prediction match the actual prescription provided by the doctor. Additionally, professional evaluations of each model are presented, ranking them using the “>” symbol (with models ranked higher indicating better performance). Results show that TCM-FTP (ShenNong) achieves strong prediction accuracy, closely matching doctors’ herb labels and dosages after fine-tuning, with clinically relevant dosage alignment. In contrast, GPT-3.5 and GPT-4.0 showed poor predictions due to a lack of specialized TCM training. This underscores the value of fine-tuning LLMs on high-quality, domain-specific data for optimal performance in specialized fields.
@@ V CONCLUSION
To deal with the lack of high-quality datasets and to improve the performance in TCM prescription predictions, we build a TCM prescription dataset DigestDS and propose TCM-FTP to fine-tune LLMs to predict herbs with the corresponding dosages. The experimental results reveal the remarkable effectiveness of TCM-FTP, surpassing previous methods by large margins in precision, recall, and F1-score. Additionally, our method achieved the best results in NMSE, effectively forming accurate herb and dosage predictions. In future work, we will continue to incorporate domain knowledge into model construction to further enhance performance, aiming to develop a practically usable prescription prediction model.
VI ACKNOWLEDGEMENT
This work is partially supported by the Natural Science Foundation of Beijing (No. L232033), the National Natural Science Foundation of China (Nos. 82374302 and U23B2062), the Fundamental Research Funds for the Central Universities (No. 2024YJS032) and Hong Kong Research Grants Council under grant (No. 16204920). We gratefully acknowledge the support of NVIDIA Corporation with the GPU computational resources used for this research.